OpenStudy (anonymous):

FIND THE MIDPOINT OF SEGMENT WITH THE GIVEN ENDPOIN. D(-2,1) AND E(-3,2)

OpenStudy (anonymous):

Find the midpoint on the x: -2.5 Midpoint on the y: 1.5 (-2.5,1.5)

OpenStudy (anonymous):

(-2-3)/2 =-5/2=-2.5 (1+2)/2=3/2=1.5